# Determination of Magnetic Symmetries by Convergent Beam Electron Diffraction

### v2

@@ V.1 LaMnAsO
The first computational example is LaMnAsO, which is an antiferromagnet with magnetic space group $P4^{\prime}/n^{\prime}m^{\prime}m$ and lattice parameters $a=b=4.114\,\mbox{\AA}$ and $c=9.0304\,\mbox{\AA}$ [45]. The magnetic moments of Mn atoms are oriented along the easy axis $c$ and are arranged in the antiferromagnetic pattern shown in Fig. 3 [45]. The symmetries of LaMnAsO that are probed by magnetic CBED are described by the point group $4^{\prime}m^{\prime}m$. We furthermore note that the point group $4^{\prime}m^{\prime}m$ is equivalent to $4^{\prime}mm^{\prime}$ obtained by rotating the unit cell through $45^{\circ}$ about $[001]$. The latter conventions have been used for the magnetic CBED groups and their mappings to magnetic point groups in Table 11.
@@ V.2 NiO
NiO is a Mott-insulating antiferromagnet below a Néel temperature of $523\,\mathrm{K}$ [46]. Its paramagnetic phase has cubic $Fm\bar{3}m$ symmetry with lattice constant $a=4.17\,\mbox{\AA}$, which becomes monoclinic $C_{c}2/c$ (BNS magnetic space group notation [47]) in the antiferromagnetic phase [48] (Fig. 5). The latter involves a doubling of the magnetic unit cell. The monoclinic axis of the magnetic unit cell points into the $[1\bar{1}0]_{c}$ direction (the cubic notation is indicated by the subscript c) and the spins are oriented along the easy axis $[11\bar{2}]_{c}$, which is perpendicular to the monoclinic axis. It is also important to note that NiO undergoes a spin-flop transition at an external magnetic field of $1.54\,\mathrm{T}$, where spins reorient perpendicular to the field direction [49]. Indeed, Loudon [22] detected the purely magnetic reflections of the spin-flopped phase by conventional broad-beam electron diffraction, where the objective lens typically imposes external fields well above the spin-flop field. The magnetic symmetry of the spin-flopped phase has not been reported yet.
@@
In contrast to LaMnAsO, we can now directly read off the symmetries from the purely magnetic reflections displayed in Figs. 6 and 7. Here, in order to display their internal structure, a gray scale was chosen that saturates the structural Bragg disks. The CBED pattern of the original phase exhibits twofold rotational symmetry as well as both $z$-reversal and time-reversal symmetries, i.e., $21_{R}1^{\prime}$ in total, see Fig. 6. Inspecting Table 8, this symmetry corresponds to the gray point group $2/m1^{\prime}$, which corresponds to the magnetic point group symmetries obtained by stripping all translations from the NiO space group $C_{c}2/c$. Indeed, the glide plane translation along the monoclinic $c$ direction does not break the symmetry for the slab oriented along monoclinic $b$-axis since $c$ is perpendicular to $b$, and leaves the slab invariant. Consequently, the glide plane symmetry translates into a simple mirror symmetry denoted by $/m$ in the $2/m1^{\prime}$ diffraction group mentioned above. Similarly, the $1^{\prime}_{c}$ symmetry (abbreviated by the subscript $c$ attached to $C$ centering in the space group symbol), i.e., the combination of time reversal and in-plane translation along $c$ corresponds to a simple time reversal symmetry $1^{\prime}$ since the translation leaves the slab invariant.
@@ VI Experimental Magnetic CBED
Having established the theoretical framework for the determination of magnetic point-group symmetries from CBED patterns, we elaborate on possible and existing experimental realizations. We identify seven main challenges and discuss their impact as well as mitigation strategies:
@@
    Signal-to-noise ratio: The previous considerations show that the modifications of CBED patterns due to magnetic scattering are small for antiferromagnets (ferromagnets will be discussed further below). Local intensity modulations in the CBED discs of LaMnAsO at $300\,\mathrm{kV}$ acceleration voltage and approximately $100\,\mathrm{nm}$ slab thickness that indicate the presence or absence of magnetic-point-group symmetries are on the order of 0.1-1% with respect to the total beam intensity. To measure such a weak modulation above the noise level, one should thus employ low-noise detectors of high dynamic range (i.e., direct electron detectors [50]) with a detection quantum efficiency close to unity, effectively reducing the noise to the inevitable Poissonian noise of the detection process. For the example of LaMnAsO, a beam current of $1\,\mathrm{nA}$ and a CBED acquisition time of $0.4\,\mathrm{s}$ yield a maximal magnetic modulation within the Bragg discs that is more than five times larger than the standard deviation of the detection (shot) noise, which is deemed sufficient for detection according to the Rose criterion [51]. If studying antiferromagnets with a doubled magnetic unit cell that give rise to purely magnetic Bragg reflections, one may furthermore focus on these reflections because they do not have a strong nonmagnetic background. However, for the example of the spin-flopped NiO, a current of $1\,\mathrm{nA}$ in combination with an acquisition time of $32\,\mathrm{s}$ is necessary to exceed a signal-to-noise level of five for measuring a modulation within the two purely magnetic Bragg discs. The larger acquisition time in this case stems from the weaker magnetic diffraction in NiO compared to LaMnAsO and the sizable thermal diffuse background due to inelastic phonon scattering. We finally also note that the relative magnitude of the magnetic signal depends on the thickness of the slab, the acceleration voltage of the electron beam, the zone-axis orientation, and the material under consideration in a complex way due to dynamic scattering, see, e.g., [44], typically allowing for some optimization of, e.g., the slab thickness, to further increase the relative magnitude of the magnetic CBED disc modulations. Considering that both beam currents in the range of $1\,\mathrm{nA}$ and acquisition times in the range of tens of seconds are easily achievable in conventional TEM, the presence or absence of magnetic-point-group symmetries in antiferromagnets should be detectable even with non-ideal detectors.
@@
    Inelastic electron scattering (excitation of phonons, plasmons, etc.) leads to a background in experimental CBED patterns, effectively lowering the signal-to-noise ratio of the magnetic Bragg signal. A significant part of this background that cannot be removed by filtering inelastically scattered electrons using an energy filter is due to inelastic phonon scattering, referred to as thermal diffuse scattering (TDS). To assess the magnitude of the TDS background, we carried out electron-diffraction simulations including TDS within the framework of the frozen phonon approximation, see Appendix D. The results show that the intensity of TDS background at room temperature in NiO at the simulated slab thicknes is approximately 10 times larger than the magnetic signal and hence does play a crucial role in experimental visibility of magnetic disks and their symmetries. To facilitate the measurement of magnetic symmetries in the purely magnetic Bragg discs with an SNR of 5 above the TDS background intensity, an acquisition time of $32$ seconds employing beam current in the range of nanoamps is required, which is experimentally well feasible. The acquisition time may be further reduced by cooling the specimen, thereby reducing the TDS background. Noting, furthermore, that the TDS background in NiO is comparatively large, due to the rather light atoms involved, the presence of the TDS background does not impede the measurement of magnetic symmetries by magnetic CBED.
@@
    The in-plane magnetic-point-group symmetries of the sample translate to CBED symmetries if the circular symmetric electron probe is well aligned with the crystallographic orientation of the slab. The accuracy of this alignment procedure is limited by the precision of the beam tilt coils of the TEM, typically allowing for fine tuning of beam tilt angles down to $100\,\mu\mathrm{rad}$ and slightly below. Thus a residual beam tilt in that range that can break point-group symmetries has to be taken into account when analyzing the presence of magnetic CBED symmetries. To asses the magnitude of such symmetry-breaking effects, we carried out CBED simulations incorporating small beam tilts with respect to crystalline zone axes, see Appendix E. The results show that small residual beam tilts result in small shifts of Bragg disks. The artificially induced symmetry breaking suppressed by generally excluding a small outer rim of the CBED discs of approximately $100\,\mu\mathrm{rad}$ in the symmetry analysis, as employed for conventional CBED [52, 53, 54].
@@
    For the determination of magnetic symmetries, the experimental realization of magnetization (i.e., time) reversal without changing the magnetic structure is helpful—but not necessary, see below. There are several methods facilitating magnetization reversal, depending on the particular magnetic material to be studied: Firstly, one can briefly heat up the sample above its Néel or Curie temperature. If there is no bias for one Néel vector orientation this would allow magnetization reversal in a randomized way. Secondly, short application of an external field may trigger a spin-flop transition in antiferromagnets, after which a random magnetization reversal may occur if no bias is present. Thirdly, one may trigger domain-wall motion by applying currents or external fields [55, 56, 57] and thereby change the magnetic domain orientation in the observed area. Note, however, that one may entirely circumvent the experimental realization of magnetization (time) reversal by recording a sufficient number of CBED patterns along different high-symmetry zone axes. For example, in the case of LaMnAsO, the magnetic point group $4^{\prime}m^{\prime}m$ may be uniquely determined from CBED patterns recorded in one or maximally two of the high-symmetry zone axes, like $\langle 100\rangle$ and $\langle 110\rangle$, see Table 11, if the underlying structural point group $4mm$ is known, e.g., from high-temperature electron diffraction or X-ray measurements.
@@
    Standard CBED and other diffraction experiments in the TEM are carried out with the sample immersed in the strong magnetic field of the objective lens of up to $2.5\,\mathrm{T}$, which permits very short camera lengths and hence recording of large CBED patterns without significant aberrations. If the influence of the magnetic field from the objective lens alters the magnetic structure under consideration, e.g., because it triggers a spin-flop transition in an antiferromagnet or reorients the magnetization in a ferromagnet, a zero-field adaption of the CBED setup is required. This may include additional instrumental modifications of the TEM, e.g., a zero-field objective lens or additional Lorentz lenses (preobjective or postobjective lenses) [58].
@@
    The $z$-reversal symmetries are generally difficult to observe experimentally because the prepared crystal slab is breaking these symmetries when the upper and lower crystal plane are not identical. As a result, one typically records CBED patterns of multiple slabs of noncolinear crystal orientations, rendering the map of diffraction symmetries to point group symmetries uniquely defined without directly probing $z-$reversal symmetries [12].
@@
    The proposed magnetic CBED can only provide magnetic point-group symmetries. The point-group transformations being part of glide plane and screw axis are typically not visible, except for some special slab orientations, where the corresponding translations are within the slab plane (see above example of spin-flopped NiO). Consequently, the detection of space-group symmetries requires additional care and may involve more sophisticated methods such as inspection of kinematically forbidden reflections or higher-order Laue zones, similar to conventional CBED [59].
@@
First experimental proofs for the feasibility of probing magnetic symmetries by CBED have already been reported for ferromagnets, where magnetic diffraction is more pronounced compared to compensated antiferromagnets. In a pioneering experimental CBED study on the uniaxial ferromagnet PrCo5, Shen and Laughlin [21] showed how the structural CBED symmetry is reduced by the presence of magnetic scattering. The structural space group of PrCo5 is $P6/mmm$, which would give rise to $6mm$ CBED symmetry along the $\left[001\right]$ and $\mathrm{2mm}$ along the $[110]$ axis, see Table 15. The experimentally observed CBED symmetry along $\left[110\right]$ is, however, clearly reduced to $m$. The theory developed above allows to explain this result. Inspecting the assignment tables for hexagonal magnetic point groups, Tables 14 and 15, two ferromagnetic magnetic point groups [60] that are compatible with $6/m$ symmetry and ferromagnetic ordering exhibit $m$ CBED symmetry. The first is $6/m$ (magnetic CBED symmetry $2_{R}mm_{R}$), while the second is $6/mm^{\prime}m^{\prime}$ (magnetic CBED symmetry $2^{\prime}mm^{\prime}1^{\prime}_{R}$). Recording a second CBED pattern along $\left[110\right]$ with magnetization direction reversed would allow to probe the presence of, e.g., the $2^{\prime}$ symmetry and hence experimental distinction between the $6/m$ and $6/mm^{\prime}m^{\prime}$ symmetries. The second experimental example pertains to CBED of La1-xSrxMnO3 by Tsuda et al. [61]. The main result is the identification of a triclinic superlattice reconstruction, presumably due to orbital ordering. Additionally, by inspecting multiple CBED patterns along different beam directions, a symmetry of $\bar{1}$ consistent with a CBED symmetry of $2_{R}$ along all axes, see Table 7, could be revealed. Indeed, $\bar{1}$ is the only triclinic ferromagnetic magnetic point group [60], necessitating no additional measurement involving magnetization reversal or different slab orientations to uniquely determine the magnetic symmetry in this case.

### v3

@@ V.1 LaMnAsO
The first computational example is LaMnAsO, which is an antiferromagnet with magnetic space group $P4^{\prime}/n^{\prime}m^{\prime}m$ and lattice parameters $a=b=4.114\,\mbox{\AA }$ and $c=9.0304\,\mbox{\AA }$ [McGuire2016]. The magnetic moments of Mn atoms are oriented along the easy axis $c$ and are arranged in the antiferromagnetic pattern shown in Fig. 3 [McGuire2016]. The symmetries of LaMnAsO that are probed by magnetic CBED are described by the point group $4^{\prime}m^{\prime}m$. We furthermore note that the point group $4^{\prime}m^{\prime}m$ is equivalent to $4^{\prime}mm^{\prime}$ obtained by rotating the unit cell through $45^{\circ}$ about $[001]$. The latter conventions have been used for the magnetic CBED groups and their mappings to magnetic point groups in Table 11.
@@ V.2 NiO
NiO is a Mott-insulating antiferromagnet below a Néel temperature of $523\,\mathrm{K}$ [Roth1958]. Its paramagnetic phase has cubic $Fm\bar{3}m$ symmetry with lattice constant $a=4.17\,\mbox{\AA }$, which becomes monoclinic $C_{c}2/c$ (BNS magnetic space group notation [belov1957shubnikov]) in the antiferromagnetic phase [Cairns1933] (Fig. 5). The latter involves a doubling of the magnetic unit cell. The monoclinic axis of the magnetic unit cell points into the $1\bar{1}0]_{c}$ direction (the cubic notation is indicated by the subscript c) and the spins are oriented along the easy axis $11\bar{2}]_{c}$, which is perpendicular to the monoclinic axis. It is also important to note that NiO undergoes a spin-flop transition at an external magnetic field of $1.54\,\mathrm{T}$, where spins reorient perpendicular to the field direction [Saito_1980]. Indeed, Loudon [Loudon2012] detected the purely magnetic reflections of the spin-flopped phase by conventional broad-beam electron diffraction, where the objective lens typically imposes external fields well above the spin-flop field. The magnetic symmetry of the spin-flopped phase has not been reported yet.
@@
In contrast to LaMnAsO, we can now directly read off the symmetries from the purely magnetic reflections displayed in Figs. 6 and 7. Here, in order to display their internal structure, a gray scale was chosen that saturates the structural Bragg disks. The CBED pattern of the original phase exhibits twofold rotational symmetry as well as both $z$-reversal and time-reversal symmetries, i.e., $21_{R}1^{\prime}$ in total, see Fig. 6. Inspecting Table 8, this symmetry corresponds to the gray point group $2/m1^{\prime}$, which corresponds to the magnetic point group symmetries obtained by stripping all translations from the NiO space group $C_{c}2/c$. Indeed, the glide plane translation along the monoclinic $c$ direction does not break the symmetry for the slab oriented along monoclinic $b$-axis since $c$ is perpendicular to $b$, and leaves the slab invariant. Consequently, the glide plane symmetry translates into a simple mirror symmetry denoted by /m/m in the $2/m1^{\prime}$ diffraction group mentioned above. Similarly, the $1^{\prime}_{c}$ symmetry (abbreviated by the subscript $c$ attached to $C$ centering in the space group symbol), i.e., the combination of time reversal and in-plane translation along $c$ corresponds to a simple time reversal symmetry $1^{\prime}$ since the translation leaves the slab invariant.
@@ VI Experimental Magnetic CBED
Having established the theoretical framework for the determination of magnetic point-group symmetries from CBED patterns, we elaborate on possible and existing experimental realizations. We identify eight main challenges and discuss their impact as well as mitigation strategies:
@@
    Signal-to-noise ratio: The previous considerations show that the modifications of CBED patterns due to magnetic scattering are small for antiferromagnets (ferromagnets will be discussed further below). Local intensity modulations in the CBED discs of LaMnAsO at $300\,\mathrm{kV}$ acceleration voltage and approximately $100\,\mathrm{nm}$ slab thickness that indicate the presence or absence of magnetic-point-group symmetries are of the order of 0.1-1% with respect to the maximal intensity in the CBED pattern. To measure such a weak modulation above the noise level, one should thus employ low-noise detectors of high dynamic range (i.e., direct electron detectors [MacLaren2020]) with a detection quantum efficiency close to unity, effectively reducing the noise to the inevitable Poissonian noise of the detection process. For the example of LaMnAsO, a beam current of $1\,\mathrm{nA}$ (that is common for CBED) and an acquisition time of $0.4\,\mathrm{s}$ yield a maximal magnetic modulation within the Bragg discs that is more than five times larger than the standard deviation of the detection (shot) noise, which is deemed sufficient for detection according to the Rose criterion [Rose1974] (lower SNR requirements are also employed frequently though). If studying antiferromagnets with a doubled magnetic unit cell that give rise to purely magnetic Bragg reflections, one may furthermore focus on these reflections because they do not have a strong nonmagnetic background. However, for the example of the spin-flopped NiO, a current of $1\,\mathrm{nA}$ in combination with an acquisition time of $32\,\mathrm{s}$ is necessary to exceed a signal-to-noise level of five for measuring a modulation within the two purely magnetic Bragg discs. The larger acquisition time in this case stems from the weaker magnetic diffraction in NiO compared to LaMnAsO and the sizable thermal diffuse background due to inelastic phonon scattering. We finally also note that the relative magnitude of the magnetic signal depends on the thickness of the slab, the acceleration voltage of the electron beam, the zone-axis orientation, and the material under consideration in a complex way due to dynamic scattering, see, e.g., [Snarski-Adamski_simulations_2023], typically allowing for some optimization of, e.g., the slab thickness, to further increase the relative magnitude of the magnetic CBED disc modulations.
@@
    Inelastic electron scattering (excitation of phonons, plasmons, etc.) leads to a background in experimental CBED patterns, effectively lowering the signal-to-noise ratio of the magnetic Bragg signal. A significant part of this background that cannot be removed by filtering inelastically scattered electrons using an energy filter is due to inelastic phonon scattering, referred to as thermal diffuse scattering (TDS). To assess the magnitude of the TDS background, we carried out electron-diffraction simulations including TDS within the framework of the frozen phonon approximation, see Appendix D. The results show that the intensity of TDS background at room temperature in NiO at the simulated slab thickness is approximately 10 times larger than the magnetic signal and hence does play a crucial role in experimental visibility of magnetic disks and their symmetries. To facilitate the measurement of magnetic symmetries in the purely magnetic Bragg discs with an SNR of 5 above the TDS background intensity, an acquisition time of $32$ seconds employing beam current in the range of nanoamps is required, which is experimentally feasible. The acquisition time may be further reduced by cooling the specimen, thereby reducing the TDS background. For instance, cooling down from room temperature to 30 K will results in reduction of required exposure times (doses) by approximately one order of magnitude, which can be seen in the following way: Measuring 10 times shorter leads to 10 times less magnetic signal, but the background goes down up to 100 times—10 times due to shorter acquisition time and additional 10 times due to 10 times reduced temperature, thus maintaining the same SNR as at 300 K at 10 times longer acquisition. Moreover, as TDS goes down, more electrons will scatter elastically instead, further increasing the SNR. Noting, furthermore, that the TDS background in NiO is comparatively large, due to the rather light atoms involved, the presence of the TDS background may have less impact on the measurement of magnetic symmetries by magnetic CBED in other materials.
@@
    Thickness gradients of the TEM lamella, strain gradients, adsorbates on the surface, as well as the beam damage may complicate the local analysis of symmetries. They typically lead to a rather diffuse diffraction intensity (background), which breaks the symmetry of CBED patterns. It is difficult to assess the impact of such effects as they crucially depend on details such as zone axis, material composition, adsorbate composition, radiation damage susceptibility, etc. A very coarse approximation may be derived by considering the intensity of the background at the position of a CBED disc and the ratio of thickness of the sample to the relative thickness increase due to a gradient, or the thickness of adsorbates layer, or the the thickness of the damaged layer, respectively. Assuming a typical thickness of the lamella for magnetic CBED of 100 nm and of the adsorbate or damaged layer of 1 nm, and assuming that the background is constantly spread out over the considered CBED regime, similar background levels as the thermal diffuse background at room temperature are obtained, posing another serious challenge for the technique. Suitable countermeasures to reduce these local symmetry breaking effects are averaging the signal over several measurement points, thereby averaging out impact of adsorbates, gradients, or damaged surface layers, careful cleaning of the sample surfaces, cooling to reduce beam damage, and careful sample preparation through focused ion beam milling, to avoid thickness gradients.
@@
    The in-plane magnetic-point-group symmetries of the sample translate to CBED symmetries if the circular symmetric electron probe is well aligned with the crystallographic orientation of the slab. The accuracy of this alignment procedure is limited by the precision of the beam tilt coils of the TEM, typically allowing for fine tuning of beam tilt angles down to $100\,\mu\mathrm{rad}$ and slightly below. Thus a residual beam tilt in that range that can break point-group symmetries has to be taken into account when analyzing the presence of magnetic CBED symmetries. To asses the magnitude of such symmetry-breaking effects, we carried out CBED simulations incorporating small beam tilts with respect to crystalline zone axes, see Appendix E. The results show that small residual beam tilts result in small shifts of Bragg disks. The artificially induced symmetry breaking may be suppressed by generally excluding a small outer rim of the CBED discs of approximately $100\,\mu\mathrm{rad}$ in the symmetry analysis, as employed for conventional CBED [LeBeau2010, Xu2018, Oberaigner2023].
@@
    For the determination of magnetic symmetries, the experimental realization of magnetization (i.e., time) reversal without changing the magnetic structure is helpful—but not necessary, see below. There are several methods facilitating magnetization reversal, depending on the particular magnetic material to be studied: Firstly, one can briefly heat up the sample above its Néel or Curie temperature. If there is no bias for one Néel vector orientation this would allow magnetization reversal in a randomized way. Secondly, short application of an external field may trigger a spin-flop transition in antiferromagnets, after which a random magnetization reversal may occur if no bias is present. Thirdly, one may trigger domain-wall motion by applying currents or external fields [Wadley_electrical_2016, Wadley18, Amin2023] and thereby change the magnetic domain orientation in the observed area. Note, however, that one may entirely circumvent the experimental realization of magnetization (time) reversal by recording a sufficient number of CBED patterns along different high-symmetry zone axes. For example, in the case of LaMnAsO, the magnetic point group $4^{\prime}m^{\prime}m$ may be uniquely determined from CBED patterns recorded in one or maximally two of the high-symmetry zone axes, like $\langle 100\rangle$ and $\langle 110\rangle$, see Table 11, if the underlying structural point group $4mm$ is known, e.g., from high-temperature electron diffraction or X-ray measurements.
@@
    Standard CBED and other diffraction experiments in the TEM are carried out with the sample immersed in the strong magnetic field of the objective lens of up to $2.5\,\mathrm{T}$, which permits very short camera lengths and hence recording of large CBED patterns without significant aberrations. If the influence of the magnetic field from the objective lens alters the magnetic structure under consideration, e.g., because it triggers a spin-flop transition in an antiferromagnet or reorients the magnetization in a ferromagnet, a zero-field adaption of the CBED setup is required. Depending on the TEM (i.e., number and strength of lenses in condenser and projective), this may be achieved by switching off the objective lens and employing the condenser and projective system of the TEM to facilitate required beam convergence and short camera lengths. If the existing optics is insufficient, additional instrumental modifications of the TEM, e.g., a zero-field objective lens or additional preobjective or postobjective lenses are suited to provide the required CBED conditions  [Shibata2019].
@@
    The $z$-reversal symmetries are generally difficult to observe experimentally because the prepared crystal slab is breaking these symmetries when the upper and lower crystal plane are not identical. As a result, one typically records CBED patterns of multiple slabs of noncolinear crystal orientations, rendering the map of diffraction symmetries to point group symmetries uniquely defined without directly probing $z$-reversal symmetries [Buxton1976].
  8. 8.
@@
    The proposed magnetic CBED can only provide magnetic point-group symmetries. The point-group transformations being part of glide plane and screw axis are typically not visible, except for some special slab orientations, where the corresponding translations are within the slab plane (see above example of spin-flopped NiO). Consequently, the detection of space-group symmetries requires additional care and may involve more sophisticated methods such as inspection of kinematically forbidden reflections or higher-order Laue zones, similar to conventional CBED [Gjonnes1965].
@@
The above list indicates that experimental realization of magnetic CBED will be very challenging. Most likely, artificial symmetry breaking by thickness gradients, adsorbates and beam damage are the biggest obstacles. Notwithstanding, the above arguments also suggest that doses and hence acquisition times required to achieve sufficient signal to noise in antiferromagnets are in reach. First experimental proofs for the feasibility of probing magnetic symmetries by CBED have already been reported for ferromagnets, where magnetic diffraction is more pronounced compared to compensated antiferromagnets. In a pioneering experimental CBED study on the uniaxial ferromagnet PrCo5, Shen and Laughlin [Shen1990] showed how the structural CBED symmetry is reduced by the presence of magnetic scattering. The structural space group of PrCo5 is $P6/mmm$, which would give rise to $6mm$ CBED symmetry along the $\left[001\right]$ and $\mathrm{2mm}$ along the $[110]$ axis, see Table 15. The experimentally observed CBED symmetry along $\left[110\right]$ is, however, clearly reduced to $m$. The theory developed above allows to explain this result. Inspecting the assignment tables for hexagonal magnetic point groups, Tables 14 and 15, two ferromagnetic magnetic point groups [Schmid1973] that are compatible with $6/m$ symmetry and ferromagnetic ordering exhibit $m$ CBED symmetry. The first is $6/m$ (magnetic CBED symmetry $2_{R}mm_{R}$), while the second is $6/mm^{\prime}m^{\prime}$ (magnetic CBED symmetry $2^{\prime}mm^{\prime}1^{\prime}_{R}$). Recording a second CBED pattern along $\left[110\right]$ with magnetization direction reversed would allow to probe the presence of, e.g., the $2^{\prime}$ symmetry and hence experimental distinction between the $6/m$ and $6/mm^{\prime}m^{\prime}$ symmetries. The second experimental example pertains to CBED of La1-xSrxMnO3 by Tsuda et al. [Tsuda2001]. The main result is the identification of a triclinic superlattice reconstruction, presumably due to orbital ordering. Additionally, by inspecting multiple CBED patterns along different beam directions, a symmetry of $\bar{1}$ consistent with a CBED symmetry of $2_{R}$ along all axes, see Table 7, could be revealed. Indeed, $\bar{1}$ is the only triclinic ferromagnetic magnetic point group [Schmid1973], necessitating no additional measurement involving magnetization reversal or different slab orientations to uniquely determine the magnetic symmetry in this case.